We installed a HV kepco power supply for the green PZT steering the YAUX beam. We did this in anticipation of the YARM alignment to take place this afternoon. We borrowed an unused DC power supply labeled "OMC Power spply" and made an appropriate SMA connection (Attachment #1), Then we set the Vset to 100.0 Volts and the current limit to 10 mA. Once we enabled the output we saw the 5.6 mA of current drawn by the eurocard in accordance with the wiki log (Attachment #2).
It may not be possible to use the PZT as per this so this work may not have a direct impact on our upcoming alignment task.
We probably bumped the ethernet cable (martian network) on c1iscey, so the FE models stopped showing up on the medm screens. When we connected it back, it seemed like the FE kept running the model and only IPC showed error. We restarted the rtcds models on c1iscey and burtrestored to today morning 5:19 am.
ETMY is properly damping now and the oplev is roughly centered as well but the OPLEV Servos are off. We did not turn them on.
We should be able to carry out our cavity arm alignment today afternoon on both arms now.